what is root x divided by x

Grade:8

1 Answers

Ashwin
9 years ago
Q – √x / X 
 
its same as you divide √2 / 2 = 1/ √2   
=> √x / x = 1/√x 
 
but if you want full derivation here it is 
since we know that √x = x ½ ------------------------------------- 1
 
and x = x------------------------------------------------------ 2
 
also am/ a  = a m-n  --------------------------------------------------3
 
from  1 , 2 and 3 we get 
√x / x =  x1/2/ x1
=> x1/2- 1  = x-1/2    
=>  √x / x = 1/√x